### v2

# Evaluation of LLMs for mathematical problem solving

Abstract
Large Language Models (LLMs) have shown impressive performance on a range of educational tasks, but are still understudied for their potential to solve mathematical problems. In this study, we compare three prominent LLMs, including GPT-4o, DeepSeek-V3, and Gemini-2.0, on three mathematics datasets of varying complexities (GSM8K, MATH500, and University datasets). We take a five-dimensional approach based on the Structured Chain-of-Thought (SCoT) framework to assess final answer correctness, step completeness, step validity, intermediate calculation accuracy, and problem comprehension. The results show that GPT-4o is the most stable and consistent in performance across all the datasets, but particularly it performs outstandingly in high-level questions of the UNSW dataset. DeepSeek-V3 is competitively strong in well-structured domains such as optimisation, but suffers from fluctuations in accuracy in statistical inference tasks. Gemini-2.0 shows strong linguistic understanding and clarity in well-structured problems but performs poorly in multi-step reasoning and symbolic logic. Our error analysis reveals particular deficits in each model: GPT-4o is at times lacking in sufficient explanation or precision; DeepSeek-V3 leaves out intermediate steps; and Gemini-2.0 is less flexible in mathematical reasoning in higher dimensions.
keywords:
Large language models, ChatGPT-4o, DeepSeek-V3, Gemini-2.0, SCoT, Mathematical problem solving
journal: .
[…]
[2]Transitional Artificial Intelligence Research Group, School of Mathematics and Statistics, UNSW Sydney, Sydney, Australia
1 Introduction
Solving mathematical problems [1] is one of the major challenge for students [2] in the current educational system. This difficulty mainly stems from the abstraction [3] and strict logical frameworks [4] that are intrinsic to mathematics, and the extremely wide field of its application areas. Students need to have a strong theoretical background and be able to apply mathematical concepts to situations in daily life. However, most students are lacking in various areas of learning mathematics, especially in conceptual understanding[5], computational anxiety[6] and the mastery of applying mathematical concepts flexibly. Mathematics requires the acquisition of advanced abstract [7] and logical thinking skills[8], as well as the mastery of a large number of formulas and theorems. However, the current education system aims at providing theoretical knowledge without sufficiently strengthening the practical skills and critical thinking[9] ability of student. Students tend to passive recipients of information with limited opportunities for active exploration and critical analysis[10]. It is difficult for students to apply their knowledge to practical problems effectively. Therefore, enriching the traditional teaching process to build up students’ practical work ability and thinking capacity is an significant issue in mathematics education.
In addition, advances in computer and network technology [11] have driven the application of educational technology (EdTech)[12] in mathematics teaching. The emergence of educational technology has transformed traditional teaching models [13]. By using varied pedagogies and active learning experiences, educational technology improves students’ learning of mathematics and their ability to problem-solve [14]. Furthermore, online learning systems [15] expose students to a wide range of resources and interactive lessons through which they can enhance their understanding of concepts and improve independent learning strategies.
Progress in natural language processing (NLP)  [16] based on artificial intelligence technologies [17] and the advent of large language models (LLMs)  [18, 19, 20] have led to a revolutionary change in educational technology, giving rise to novel possibilities for mathematics education. The application of LLMs  [21, 22, 23] has fundamentally changed the design of human-computer interaction in educational settings  [24, 25] and initiated a revolutionary change in instructional approaches. In contrast to conventional systems that rely on predefined rules, large language models (LLMs) demonstrate more advanced language comprehension and generation capabilities  [26]. Not only do they accurately comprehend user intent and produce natural, fluent text content  [19], but they also adapt their output dynamically to feedback, thus creating a more semantically rich and interactive learning experience  [27]. These characteristics create new opportunities for applying LLMs and it has significantly driven the evolution of human-machine collaboration models in multiple fields  [28], including educational tutoring, customer service, programming assistance, and interactive communication support, enabling them to achieve better knowledge acquisition and problem-solving in learning, work. Large language models have demonstrated significant potential in mathematics education. Students have widely utilised LLMs to obtain problem-solving [18] strategies and answers [18], significantly shortening the time required to brainstorm and solve problems. However, whether this has truly improved learning outcomes remains to be explored [29]. Additionally, LLMs have the ability to accurately analyse the student’s learning situations[30] and dynamically adjust the presentation of content based on their pace and preferences[30], which can help promote "anytime, anywhere learning" teaching methods[31]. As LLM technology continues to advance, the potential for its use in education is becoming more widespread. The deep integration of mathematics education with artificial intelligence [32] is opening up new learning pathways and driving the construction of efficient, personalised learning platforms, thereby improving teaching quality and efficiency and promoting the comprehensive development of students[33]. Modern LLMs are typically built on deep learning models [34, 35] and widely applied in NLP tasks [36][37]. These models demonstrate strong capabilities in text understanding and generation capabilities and are widely applied in machine translation, text summarisation, and code generation [38]. Their task versatility stems from large-scale pretraining [39] and their capacity to manage tasks in both zero-shot and few-shot learning [40]. LLMs include GPT-4o [41], Gemini-2.0[42], and DeepSeek-V3[43] capable of text generation, logic reasoning, and multilingual processing[44]. Despite their advantages, LLMs encounter several challenges, such as hallucination (i.e., generating false information[21], and inherent data biases reflecting social and cultural stereotypes [45][46][47][48]. Although current LLMs demonstrate outstanding capabilities in mathematical problem solving, formula derivation, and theorem explanation [49], they remain weak in symbolic computation, multi-step algorithms and formal theorem proving [50]. To address these shortcomings, researchers have developed tools such as Wolfram Alpha[51] to enhance computational accuracy and logic rigour, and have improved programming performance through models like GitHub Copilot[52], GPT, and Code Llama[53]. However, these models still face challenges such as a lack of logical consistency[54], difficulty handling large-scale dependency structures, and generating errors or fictional functions when dealing with complex tasks. Algorithmic breakthroughs are still required to meet the demands of complex application scenarios [55], and how to efficiently utilise LLMs to improve learning outcomes while ensuring fairness [56] has become a focal point of current academic research.
Therefore, it is necessary to conduct systematic evaluations of LLMs using mathematical benchmark [57] datasets such as GSM8k [58] and MATH, which are widely used in mathematical tasks. And in order to study the performance of LLMs in advanced-level mathematical problems, we plan to construct the University of New South Sales problem dataset, which is collected by using assignment and final exams from the Mathematics and Statistics School. These evaluations help us to identify the strengths and weakness of models across different tasks [59], and provide empirical support for model optimisation. Although existing studies have fully demonstrated the ability of LLMs in mathematical and programming tasks, their shortcomings in logical reasoning and computational accuracy[60] still limit their widespread application. Therefore, it is imperative to continue optimising LLMs from the perspectives of algorithm design and technical pathways to enhance their overall performance and meet the increasingly complex and diverse demands of practical applications[61].
Our study is based on this premise and aims to evaluate the performance of three mainstream LLMs, GPT-4o, DeepSeek-V3, and Gemini-2.0, in solving mathematical problems. We will examine their capabilities in terms of accuracy, reasoning quality, completeness, and consistency using a set of problems covering different levels of difficulty. We will use an automated quantitative evaluation framework to provide theoretical support and practical evidence for their application in education.
Our study will focus on the following contributions:
  • 1.
    Systematically evaluate the accuracy and efficiency of diverse LLMs in solving mathematical problems with different levels of complexity.
  • 2.
    Compare the strengths and weaknesses of each model in terms of mathematical reasoning, accuracy and correctness of solutions, and problem-solving adaptability.
  • 3.
    Provide advice about different types of LLMs for mathematical problems that can be used to develop innovative approaches in the development of assignments and exams in mathematics education in secondary schools and universities.
2 Related Work and Background
As the application of large language models in mathematics education continues to expand, questions regarding their capabilities, impacts, and suitability have increasingly drawn academia attention. As previously mentioned that LLMs have not only transformed the way humans interact with machines but also provided new tools and approaches for solving mathematical problems. However, their performance in actual educational settings has also been met with controversy and challenges. This section aims to review relevant research, starting from the issue of academic integrity, to conduct an in-depth analysis of the technological evolution of LLMs and their performance in mathematical tasks, thereby providing theoretical support for future empirical evaluations.
2.1 Academic Integrity and the Boundaries of AI Use
Although LLMs demonstrate outstanding performance in terms of generative capabilities and problem-solving efficiency, their tendency to "directly provide answers" has raised concerns about academic integrity[62]. In an academic environment, students may bypass the thinking process and rely directly on models to complete assignments or exams. This phenomenon not only undermines students’ self-directed learning but also may lead to academic misconduct[63]. In addition, due to the easy accessibility of LLMs, the risk of their misuse for online cheating in examination scenarios increases significantly, which seriously threatens the validity of traditional assessment methods[64]. Therefore, how to balance the use of technology with learning outcomes has become a focal point of current discussions.[65]. At the same time, in recent years, some researches on the capabilities of LLMs in mathematical and programming capabilities is promoted with the initial aim of improving the quality of computer-assisted learning[66]. As an educational tool, LLMs can personalise content, provide real-time feedback, and accelerate the learning process. In the field of scientific research, LLMs can reduce the burden of researchers in routine tasks, so that they can devote more energy to high-level innovative work[41]. The future development direction should focus on improving the explanatory ability and logical reasoning level of the model, as well as improving the evaluation system and anti-cheating mechanism to ensure the suitable and responsible application of artificial intelligence in education[67][68].
2.2 Model Selection and Technical Foundations
The development of large language models is based on breakthroughs in deep learning [69] technology. In particular, neural network  [70, 71] models with the Transformer architecture at their core have enabled language processing systems to model complex contextual relationships  [72, 73] more efficiently. Compared to traditional machine learning [74] methods that rely on manually designed features [75], such as linear regression and support vector machines [76], deep learning uses multi-layer network structures to automatically extract multi-level semantic representations from large-scale data, significantly improving the generalisation and task adaptability of models.
In the field of natural language processing (NLP), the introduction of the Transformer architecture [77] marked the entry of generative language models into practical use [78]. Through large-scale pre-training language modelling tasks, models can learn language structures, semantic relationships, and even a certain degree of reasoning ability. Subsequently, combined with supervised fine-tuning [79] and human feedback optimisation mechanisms (such as RLHF), large language models began to acquire the ability to execute complex instructions and generate multi-step reasoning processes[80]. In recent years, various models such as the GPT series, DeepSeek system, and Gemini framework have been released based on this technical framework and have achieved significant results in tasks such as language understanding, code generation, and mathematical modelling [34].
Our study selected three representative models-GPT-4o, DeepSeek-V3, and Gemini-2.0-as the subjects for evaluating mathematical problem-solving capabilities. The following sections will introduce their technical features and task performance, providing a foundation for subsequent method and result analyses.
2.2.1 ChatGPT and GPT-4o
GPT developed by Open AI, which is an LLM that can solve mathematical problems[81, 78] based on the Transformer model[82]. It mainly introduces the self-attention mechanism[83] to achieve a global modelling of word relationships in the input text and then profoundly understands and generates the context. Therefore, the model can understand the structure of mathematical expressions, which helps solve complex mathematical problems. GPT-based LLMs have been pre-trained using a self-supervised learning objective [84] on large-scale text corpora, enabling them to acquire general language understanding and preliminary reasoning capabilities. Reinforcement Learning from Human Feedback (RLHF) [85] has been used to fine-tune the model through instruction-based learning, enabling it to generate outputs that better align with human preferences in language style and logical expectations when solving mathematical problems. GPT has performed well in NLP [86] and can efficiently complete various complex tasks, such as dialogue generation, text writing, logical reasoning, and programming assistance [87]. It provides efficient help at the exploratory, heuristic, and teaching levels, and it is a powerful auxiliary tool for research, teaching, and mathematics learning.
Multiple studies have indicated that GPT demonstrates a certain degree of generalisability in mathematical tasks, particularly in problems with clear structures [88, 89, 90, 91]. It can handle basic computational and symbolic reasoning problems, such as algebraic operations, equation solving, function analysis, and limit calculations, and typically provides clear step-by-step derivations. In modelling and application-oriented tasks, GPT also demonstrates some auxiliary capabilities, assisting in the construction of mathematical models for statistical, probabilistic, and optimisation problems, and attempting to explain the underlying principles involved. In programming and numerical computation, the model can generate basic code in languages such as Python, Matlab, Stata, and R based on task requirements, supporting numerical simulation and data analysis. Notably, GPT demonstrates strong understanding capabilities when converting natural language problems into mathematical expressions, providing structured problem-solving pathways. Additionally, it is suitable foe explaining complex mathematical concepts such as linear algebra, differential equations, and number theory, thereby enhancing learners’ understanding of mathematical principles.
2.2.2 DeepSeek
DeepSeek is a general-purpose large language model (LLM) independently developed in China [92] , designed to support bilingual understanding, logical reasoning, code generation, and mathematical modelling. Since its initial release in 2023, the DeepSeek team has iteratively advanced the model architecture, successively launching DeepSeek-V1, V2,and V3  [93], in an effore to establish a scalable, controllable, and continually evolving Chinese LLM framework [94].
The earliest version, DeepSeek-V1 [94] followed the standard Transformer architecture and demonstrated excellent natural language instruction-following capabilities [94]. It could handle basic arithmetic operations, text problem analysis, and algebraic equation solving tasks. However, its performance in multi-step reasoning tasks was limited, with frequent inconsistencies and a lack of support for symbolic computation or tool-based workflows[95, 43].
DeepSeek-V2[94] introduced the Mixture of Experts (MoE) architecture [96, 97], which improves inference efficiency while maintaining parameter scale, and enhanced its ability to handle math and programming tasks through the sub-model DeepSeek-MATH [98]. This version begins to support Chain-of Thought (step-by-step reasoning) [99] output and can invoke Python/SymPy [100] and other tools to complete symbolic computation. It demonstrates good performance in solving mid-to-high-level mathematical problems on the GSM8K, MATH, and SVAMP [101] datasets, particularly in tasks involving functions, geometry, and algebra.
DeepSeek-V3[94] returns to a fully parameterised dense architecture to further improve language logical consistency, cross-task generalisation, and accuracy in Chinese and English expression [102]. This version shows significantly enhanced stability in multi-round question-answering, scientific writing, and complex mathematical problem handling, with improved text understanding and ambiguous problem resolution capabilities [103]. Additionally, DeepSeek-V3 demonstrates strong explanatory reasoning capabilities, clearly presenting reasoning chains, and has achieved performance comparable to GPT-4-Turbo in code invocation and formula generation. It shows good consistency in a bilingual environment and can collaborate with subsystems such as DeepSeek-Coder [104] and DeepSeek-VL [93] to support mixed-media mathematical tasks (e.g., chart questions and geometric recognition).
Overall, the DeepSeek model demonstrates continuous evolution in mathematical language modelling, code generation, and symbolic reasoning. Its latest version has laid the foundation for participating in complex reasoning tasks, though there is room for improvement in flexibly addressing open-ended questions and adjusting strategies.
2.2.3 Gemini
Gemini is an LLM developed by Google DeepMind[105], optimised for scientific computing and multi-model reasoning scenarios. It demonstrates strong capabilities in science and engineering,including mathematical problem solving  [106]. Gemini integrates logical reasoning, symbolic computation, and code generation modules, supporting a wide range of applications from basic question types to complex scientific research problems. In comparison to other mainstream LLMs, Gemini places greater emphasis on the completeness of the reasoning chain and process interpretability [107].
In terms of mathematical code generation, Gemini also performs stably, supporting mathematical libraries such as Python, NumPy, and SymPy [108], and capable of completing tasks such as function plotting, symbolic solving, and numerical simulation. Its integrated "language-formula-program" [109] output method facilitates the complete closed loop from natural language questioning to mathematical modelling to program verification, making it particularly suitable for automated preliminary modelling of scientific research problems.
3 Methodology
We first present the framework used to evaluate selected LLMs for mathematical problem-solving. Our approach combines multiple benchmark datasets along with University assignments and exams. We incorporate controlled prompting strategies, with both quantitative and qualitative evaluation.
We need to conduct a series of controlled experiments using standardised mathematical problems to evaluate the mathematical problem-solving capabilities of LLMs. Therefore, we select three mainstream benchmark datasets- GSM8K, MATH, and the problem sets of the University of New South Wales [110] to cover different difficulty levels and problem types. In addition, we explored the impact of different prompting strategies, including zero-shot, few-shot, and Chain-of-Thought (CoT)[111], to examine each model’s ability to perform multi-step reasoning.
3.1 Mathematical problems
To ensure representativeness and reliability of the evaluation results, this study selected three sets of mathematical problems from different levels and backgrounds: GSM8K, MATH 500, and the graduate course question bank from a selected University. These sets cover a wide range of problem types, from basic arithmetic and algebraic operations to advanced statistical reasoning and optimisation, and are characterised by diverse structures and distinct language styles.
  • 1.
    GSM8K: This dataset was released by OpenAI, which contains 8,500 application-based problems at the elementary and middle school levels. The language is colloquial, and the solution paths typically range from multiple steps, making it suitable for testing a model’s ability to process basic mathematical language and logic.
  • 2.
    MATH500: This dataset has 500 different difficulty and domain questions selected from the MATH benchmark, covering algebra, sequences, geometry, functions, and other fields. The question stems are close to academic teaching materials, with long chains of reasoning, suitable for analysing multi-step thinking abilities.
  • 3.
    University Graduate Question Bank: This dataset consists of exam and assignment questions covering probability inference, matrix decomposition, optimisation theory, and other courses. The language is rigorous, and the expression style is academic. Some questions have open structures, conceptual explanations, and non-standard expressions. The details about this dataset can be shown in Table LABEL:tab:data_unsw.
Table 1: Research data sources
Dataset Size Level Source
GSM8K 5000 Primary OpenAI public dataset
MATH 500 500 Intermediate Hendrycks et al.(2021)
University Problem 130
All questions have undergone standardised pre-processing, including English standardisation, formatting, symbol conversion, and removal of non-structural redundancy. Additionally, we manually labelled each dataset with its knowledge domain (e.g., functions, linear algebra), problem-solving type (computational, reasoning, conceptual), and difficulty level (primary, intermediate, advanced) to support subsequent hierarchical analysis. We summarise the source, quantity, and difficulty level of each dataset category in Table 1.
We balanced the types and difficulty levels of questions during sampling to avoid assessment bias due to an imbalance in question types. For example, in MATH500, we avoided selecting questions from several categories while ensuring that all disciplines (such as statistics, optimisation, and financial mathematics) were covered in the dataset. In addition to supporting automatic comparison of model output results, we have manually annotated the standard answers and mathematical expression structures for each question. For some questions, we have established tolerance rules (e.g., accepting equivalent expressions and allowing a certain margin of error for floating-point numbers.
3.2 Automated LLM Evaluation Framework
Our evaluation focuses on the main performance dimensions: accuracy of final answers, correctness of intermediate computations, and symbolic processing capabilities. We combine quantitative metrics with qualitative error analysis to gain a comprehensive understanding of the problem, classify common failure patterns, and highlight the advantages and limitations of each model.
In particular, we expand the existing benchmark by introducing a set of original high-level problems from the University of New South Wales curriculum. These problems cover multiple topics and require varying degrees of abstract thinking, logical reasoning, and practical application skills, thereby providing a more rigorous test of mathematical ability.
Although previous studies have examined the mathematical performance of LLMs, such as mathematical reasoning, researchers have limited their evaluations to single models or narrow problem domains assessments [112]. In contrast, our study provides a broader cross-model comparison across multiple datasets and evaluation strategies. We aim to provide more practical insights into the current capabilities and limitations of LLMs in mathematics education by incorporating international benchmarks and university-level datasets with the following details.
  • 1.
    GSM8K [113] (Focuses on basic mathematical reasoning);
  • 2.
    MATH 500[114] (Contains medium to high difficulty questions);
  • 3.
    University Problem dataset (real academic assessment scenarios (Table LABEL:tab:data_unsw))
We designed an automated evaluation framework (Figure 1) to achieve reproducible evaluation of multiple LLMs on thousands of mathematical problems. The overall design emphasises systematic and practical approaches, aiming to construct a fair, transparent, and reproducible evaluation framework[60] to get a better understanding of the actual performance of different LLMs in solving mathematical problems  [^1] https://docs.google.com/presentation/d/1CJbzXxncH7OPakhvoCSlXwrQ0s5JOyt7N3s_SitkQ40/edit?usp=sharing.
Figure 1: Automated framework for evaluation of LLMs in mathematical problem-solving.
  • 1.
    Stage 1 - Problem definition: We establish the foundations of the evaluation using the SCoT (Scoring Chain-of-Thought (SCoT) [115] framework that emphasises systematic evaluation of the completeness and logical consistency of the reasoning chain to test the performance of LLMs, which serves as the basis for the four subsequent evaluation indicators design.
  • 2.
    Stage 2 - Dataset pre-processing: We select many mathematical questions from the GSM8K, MATH500, and University datasets (Table 1), covering multiple levels from basic arithmetic to advanced reasoning. After pre-processing, we classify the questions according to knowledge points and difficulty levels to facilitate subsequent layered evaluation and error analysis. We need to balance question types to ensure that different kinds of problems are represented in the evaluation.
  • 3.
    Stage 3 - LLM Invocation: We invoke three language models- GPT-4o, DeepSeek-V3, and Gemini-2.0 via API interfaces to answer the given questions. We use a consistent prompt structure and prompt each question multiple times to verify the consistency of the model outputs. Furthermore, we verify the original model responses, timestamps, and other metadata as the basis for subsequent quantitative analysis and error classification,as shown in Table 2.
  • 4.
    Stage 4 -Result Evaluation: We review the model output results using error metrics from the literature [116] as shown in Table 3.
    Table 2: The Methods of the Evaluation for LLMs
    Category Objection Description
    Qualitative Analysis Difficulty Level The aim is to study the performance of large language models in solving mathematical problems of different difficulty levels within the same discipline.
    Domain(Subject) The aim is to analyze the performance of large language models in solving mathematical problems under the same difficulty level across different disciplines.
    Quantitative Analysis GSM8K A quantitative assessment of 5000 questions from GSM8K was conducted, with a focus on the correctness of the final answers and the rationality of the steps.
    MATH500 A quantitative assessment was conducted on 500 questions selected from the MATH benchmark test to verify the performance of the language model in the multi-step reasoning process.
    Control Variables Problem Type Types include proof-based, conceptual, procedural, and word problems.
    Difficulty Level All the questions were manually classified into 5 levels based on the length of the questionnaires. This classification controlled the variable of task difficulty, making the model more scientific when comparing performance horizontally.
    Table 3: Evaluation dimension description based on the SCoT framework.
    Dimension Explanation
    Accuracy Check if the output answer is consistent with the correct answer, accepting floating point errors and equivalent expressions
    Reasoning Quality Check if the explanation process is logically clear and the steps are reasonable
    Completeness Check if all sub-questions can be answered based on both explicit and implicit conditions when generating results, and whether there are any skipped steps
    Consistency Check if the model outputs consistent results for the same question across multiple runs
    Accuracy of Intermediate Calculations Check if each intermediate value is correct and free of calculation or clerical errors
    Understanding of the Question Check if the model correctly interprets key conditions and understands what is being asked
    These evaluation dimensions refer mainly to the principles proposed by the SCoT framework and combine the types of problems we observed in our error analysis. Based on the automatic judgment of whether the model answers are correct, we manually analysed the model outputs for some representative questions to identify common failure patterns and logical defects [117]. We will brainstorm and reach a consensus on boundaries or uncertain answers to avoid being more subjective in grading. Qualitative analysis[118] was considered to observe the answering style, language expression, and development of ideas of different models. This can help us to gain an in-depth understanding and horizontal comparison of model behaviour.
3.3 Evaluation Metrics and LLM Prompting Strategies
The LLM evaluation focuses on two aspects: algebraic expression computation and mathematical reasoning problems [119]. Since LLMs are generative models, ensuring that the generated answers are accurate becomes very tricky, especially for mathematical reasoning tasks [120]. In the case of computational problems, we prefer to provide standard reference answers so that we can easily compare the model-generated answers to the correct answers. This facilitates the assessment of the model’s performance in terms of numerical accuracy. We adopt the automatic proof verification method for the mathematical proof problem. Instead of relying on manual evaluations, we test the performance of the model by establishing the SCoT framework and an automatic scoring method inspired by recent advances in LLM-based reasoning verification [ren2023self]. It provides a more objective criterion for evaluating mathematical reasoning problems and improves the efficiency of the evaluation process [122]. We also provide a prompting mechanism to assess the ability to correct their mistakes. If a model provided an incorrect answer, it would be prompted again to solve the problem, and the corrected answer would be evaluated.
Our framework covers the whole process from initial setup, data loading, numerical processing, model execution, result comparison, to error recording and statistical output (Figure 2). Each part is designed as an independent and interconnected unit, which is easy to expand and tune in different scenarios and provides an effective means for teaching, testing, model performance comparison and prompt tuning.
Figure 2: Workflow of evaluating LLMs for mathematical problem datasets.
  • 1.
    Step 1 - Initialisation: In the framework’s initial stage, we should first complete the necessary configuration work to ensure that all parameters and modules are ready. Researchers need to specify the model name used (such as gpt-4o, gpt-3.5-turbo), then they should set the API Key[123], and specify the evaluation data file path (such as train.jsonl) and the number of samples (such as 500 questions). This part ensures the smoothness of data calls and model access in subsequent experimental processes.
  • 2.
    Step 2 - Loading Data: Our framework considers the problem (dataset) from a JSONL [124] file to prepare the LLM input. We define each row as a JSONL object that contains the description of the math problem and the standard answer information [125]. Parsing this JSONL data line by line builds a complete list of questions. It provides a standardised data source for subsequent model inputs, ensuring that each sample is processed and evaluated in the desired format [126].
  • 3.
    Step 3 - Numerical processing and data cleaning: We must perform uniform numerical treatment on the model output and the standard answer[127] to make the comparison of answers more accurate. The core operations of this stage include removing non-numeric characters from answers, formatting the data in the form of "1,200%" and converting it into the standard number "1200"; converting integer form to string representation (for example, converting 1200.0 to "1200"). At the same time, we also need to use regular expressions to extract the last digit part of the answer. The primary purpose of this step is to eliminate comparison errors due to format differences, thus ensuring the accuracy of subsequent evaluation results[128].
  • 4.
    Step 4 - LLM execution: After complete data pre-processing, the framework automatically sends a question request to the LLMs. The system prompt word is pre-set in the request to ask the model to return only a "pure number" answer without any explanation [128] to ensure that the LLMs output results are in a uniform format. This step not only effectively controls the content of the answer output so that it can only output what we want, but also allows the evaluation process to focus only on numerical comparison. It reduces the risk of ambiguity in the model’s understanding of the question  citelaskar2024systematic.
  • 5.
    Step 5 - Evaluate LLM solution: Our framework needs to clean the LLM output and the standard answer numerically, and then perform strict comparison in the form of a string matching [91]. In this way, it can accurately judge whether the solution of the model is consistent with the standard answer, to obtain the correct result of each question, and calculate the accuracy rate[60]. This process is the core of assessing the mathematical ability of LLMs and provides direct data support for mathematicians [129].
  • 6.
    Step 6 - Qualitative Error Analysis: Based on the evaluation results, representative model outputs are manually analysed to identify typical errors in the reasoning chain. Typical errors in the reasoning chain are identified. This stage supplements the details that cannot be captured by the automated scoring and enhances the explanatory nature of the assessment.
  • 7.
    Step 7 - Log Error Cases: The evaluation system records all cases of errors [130] to further analyse the model’s performance on a specific topic. The record’s content includes the question number, the question, the standard answer and the output of the LLMs. Therefore, we can further explore the causes of errors, and provide data and empirical support for model improvement and prompt optimisation
  • 8.
    Step 8 - Summary statistics of results: Following the above evaluations, we need to count and print the overall evaluation results. The system will compute the total accuracy based on the comparison results. Meanwhile, the evaluation framework will also output the evaluation summary as a JSONL file or print it out directly. It is human-readable material and a reference for immediate judgment of model performance.
3.4 Evaluation
We utilise our LLMs evaluation framework to evaluate the performance of the GPT-4o, DeepSeek, and Gemini-2.0 models in providing accurate responses. The evaluation framework covers four main objectives to measure the accuracy and reliability of the model comprehensively.
  • 1.
    Accuracy: This will help us determine whether the answer finally output by the LLM is consistent with the standard answer. It is the most direct indicator to measure the model’s problem-solving capability. The accuracy of the solution has been commonly used as the primary evaluation criterion of LLMs [106].
  • 2.
    Reasoning: Refers to the logical clarity, rationality and degree of conformity to the semantics of the question of the intermediate reasoning steps generated by the model. Reasoning quality focuses on whether there are logical problems, such as skipping steps, circular reasoning, and unfounded assumptions [131].
  • 3.
    Completeness: Even if the final answer is correct, if the intermediate steps are missing or part of the reasoning is skipped, it is considered lacking in completeness. Therefore, completeness ensures whether the LLM covers all the key steps required to solve the problem[132].
  • 4.
    Consistency: This enables us to evaluate if the LLM can ensure consistency in logical structure[133], and answer output while avoiding unstable reasoning through randomness or language fluctuation.
4 Results
We comprehensively evaluate the advantages and disadvantages of LLMs in mathematical problem-solving tasks by examining the key indicators, including accuracy, reasoning quality, completeness and consistency. Furthermore, the evaluation framework utilises the Chain of Thought [134], Few-sample and Zero-sample [135] techniques, and multi-round verification processes [136] to improve the accuracy of the evaluation results.
4.1 LLMs for solving GSM8K problems
Table 4: Accuracy (%) of LLMs for GSM8K Datasets
Model GPT-4o DeepSeek  Gemini
Accuracy 50.0% 42.6% 53.9%
We systematically evaluate the selected LLMs based on the GSM8K dataset, denoting basic mathematical reasoning tasks, as this dataset comprises 5,000 questions. Although its difficulty level belongs to elementary mathematics, covering typical question types such as multi-step problems, basic arithmetic operations, and unit conversions, the questions are relatively simple regarding knowledge difficulty [50]. However, this dataset poses a comprehensive challenge to LLMs’ logical operations, organisation and computational capabilities.
We investigated more about the errors contributing to this mediocre result, including interruptions in the reasoning chain, omissions of steps, misunderstandings of quantitative relationships, and errors in unit conversion. In the case of GSM8K, all three models sometimes arrives at the correct answer. However, they cannot generate intermediate steps due to an unstable reasoning process or improper handling of details, which reflects their lack of coherence and robustness in multi-step tasks, as shown in Figure LABEL:fig:sample_question_answer.
Although all three models can assist in solving basic maths tasks, they can still fail to consistently provide stable and plausible correct answers when faced with reasoning-based problems such as GSM8K. This suggests that the reasoning capabilities of LLMs cannot yet meet the practical needs for robustness and interpretability of fundamental mathematics in an educational setting.
Table 5: Common reasoning errors in GSM8K (sample-based)
Error Type Description
Step Skipping Model omits key intermediate steps required to reach the final answer.
Unit Confusion Incorrect unit conversions.
Misinterpretation Misunderstand conception and comparative or conditional phrases.
Arithmetic Mistakes All steps are correct but the final answer is wrong.
Table 4 shows that Gemini performs relatively best on the GSM8K dataset, with an accuracy of 53.9%. In addition, it demonstrates a more complete reasoning path compared to GPT-4o (50.0%) and DeepSeek-V3 (42.6%)  [^2] https://github.com/sydney-machine-learning/LLM-mathematics. To better understand this difference, we qualitatively (manually) review representative model responses, as summarised in Table 5. Gemini produced fewer instances of step skipping, unit confusion, and misinterpretation of the question intent. In contrast, we also found that DeepSeek’s reasoning chain is the most unstable, with frequent errors in unit conversion and logical expression, and is at the bottom of the three in terms of stability and accuracy.
4.2 LLMs for MATH500
Table 6: Accuracy (%) of LLMs for MATH500 Datasets
Model GPT-4o DeepSeek  Gemini
Accuracy 36.8% 51.1% 54.7%
Figure 3: Accuracy heatmaps of GPT-4o, Gemini-2.0, and DeepSeek-V3 on MATH500 by subject and difficulty level.
In the case of MATH500 problems using LLMs, we evaluate the accuracy of the answers in terms of two dimensions: the difficulty level of the questions and the domain involved in the questions. MATH500 consists of problems from seven subjects: Algebra, Counting and Probability, Geometry, Intermediate Algebra, Number Theory, Prealgebra, and Precalculus. The issues in Math500 are more theoretical than those in GSM8K and are mainly used to test the application of knowledge and formulas. We subdivided the difficulty level into five levels based on the length and characteristics of the words in the questions, and the specific division criteria are outlined in Table 7.
Table 7: MATH500 - difficulty level
Level Features Average Length Sample
1 Single-step arithmetic, explicit expression, no transformation required 20-25 words Solving quadratic equations, simple substitution
2 Multi-step but clearly structured, no hidden conditions 25-30 words Area calculations, unit conversion problems
3 Abstract expressions, multi-step reasoning, implicit relationships 30-35 words Quadratic properties, categorical reasoning
4 Multi-step reasoning, need for constructive method, non-intuitive structures 35-40 words Number theory analysis, path counting problems
5 Highly abstract, complex conditions, proof-like problem structures >40 words Conditional probabilistic chained derivations, function constructions
Table 8: Accuracy (%) of LLMs on MATH500 by Difficulty Level and Subject
Category Item DeepSeek Gemini GPT
Level Level 1 58% 79% 62%
Level 2 51% 66% 63%
Level 3 44% 61% 45%
Level 4 36% 53% 44%
Level 5 29% 45% 26%
Subject Prealgebra 52% 69% 51%
Count/Probability 37% 45% 45%
Algebra 40% 49% 35%
Geometry 34% 53% 30%
Number Theory 31% 48% 35%
Precalculus 30% 50% 26%
Intermediate Algebra 38% 38% 25%
We present the performance of the respective LLMs for the MATH500 dataset in Table 8. Although it handles lower and intermediate difficulty problems with reasonable success, GPT-4o’s accuracy declines significantly on higher-level tasks. GPT-4o achieves an accuracy of 62% on Level 1 and 63% on Level 2 questions, comparable to DeepSeek but lower than Gemini’s performance (79% and 66%, respectively). These results indicate that GPT-4o can handle basic computation and well-structured problems with relatively stable reasoning. Gemini 2.0 performs relatively consistently in questions of lower difficulty (level 1 and level 2), reaching an average accuracy of about 79% and 65%, respectively. However, when the difficulty level is raised to level 3 and above, the accuracy rate drops significantly and gradually decreases with increasing difficulty. Gemini 2.0 reports the lowest accuracy rate in level 5 questions, achieving only about 35%. This trend reveals Gemini’s weak ability to deal with complex mathematical tasks such as abstract algebra and combinatorial counting, and in particular, cannot master multi-step logical reasoning paths. We can see that the overall accuracy of DeepSeek is not high; level 1 has the highest accuracy rate of 62.8%. However, as the difficulty increases, the accuracy rate rapidly decreases from level 2 and is all below 50% and is less than 30% afterwards. This indicates that DeepSeek does not perform well in handling some mathematical problems, especially those that are slightly more complex or difficult.
Table 8 demonstrates that Gemini performed relatively well in Prealgebra and Geometry, with accuracy rates of approximately 70% and 54%, respectively. However, Gemini’s answer accuracy is the lowest at 38% and 45% on the Intermediate Algebra and Counting and Probability questions. This result further reveals that the model has better generalisation ability in simple questions with clear structure and rules. Still, it shows some limitations in compound questions involving implicit reasoning and conditional construction. Among the subjects covered by DeepSeek, only Prealgebra has an accuracy rate of exactly 50%, while the other subjects have an accuracy rate of around 30%. Among them, the worst-performing subject is Intermediate Algebra, with an accuracy rate of only 18.6%, indicating that DeepSeek does not perform well in dealing with mathematical problems that do not have a complete logical framework and conditional limitations.
In summary, GPT-4o performs reasonably well on low to moderately difficult problems, but its accuracy drops significantly on complex issues requiring multi-step reasoning. Despite its excellent language generation capabilities, the model often suffers from incomplete reasoning chains, skipped steps, or logic breaks in real-world mathematical tasks, especially in scenarios where the response length or reasoning time is limited. In contrast, DeepSeek-V3 has the lowest overall accuracy, and its performance decreases the most as the difficulty of the questions increases. The model occasionally performs well on basic arithmetic or probability questions, but does not perform well on problems with dense algebraic structure or open logic. Its output often lacks key intermediate steps, and the reasoning process lacks coherence. Gemini-2.0 demonstrates the most stable and scalable mathematical reasoning on the MATH500 dataset. GPT-4o achieves a balance between language and problem-solving ability in easy tasks. DeepSeek-V3 has a weaker overall performance and is mainly suited for low-difficulty tasks or scenarios that require tool-call support. In future evaluations, using Chain-of-Thought (CoT) prompting [137]can help by encouraging the model to show its step-by-step thinking. CoT can improve both performance and explainability, giving a more accurate picture of what the model can do.
To investigate the interaction effect of the model on different branches of mathematics and the difficulty of the questions, this study constructs a heat map of the accuracy of Gemini-2.0 on the MATH500 dataset based on the dimensions of "subject area" and "difficulty level", as shown in Figure 3. The heat map shows that questions in Precalculus and Number Theory are primarily distributed in the lower difficulty levels, with almost no problems in level 4 and level 5 involved. In contrast, problems in pre-algebra are widely distributed in levels 1 and 3, and some appear in level 2 and level 5, indicating that the accuracy in this area contains both basic introductory questions and more complex extensions. The heatmap (Figure 3) in the case of DeepSeek on the MATH500 dataset shows that most of the questions of Precalculus, Prealgebra, Number Theory, and Intermediate Algebra have relatively low difficulty levels, mainly distributed at levels 1 and 2, with almost no high difficulty questions. In contrast, most Geometry problems are distributed around level 3, probably due to their higher complexity and the need for more constraints. Therefore, DeepSeek has a few advantages in solving simple problems, such as basic pre-algebra or counting and probability. Although DeepSeek performs averagely in other aspects, the high difficulty problems it solves demonstrate its potential for improvement.
5 Conclusion
This study comprehensively evaluated the performance of GPT-4o, DeepSeek-V3, and Gemini-2.0 in mathematical problem-solving tasks in five dimensions, covering three types of datasets, including basic (GSM8K), intermediate (MATH500), and advanced (University). This study presents an automatic scoring framework to evaluate LLMs on mathematical problems featuring multidimensional qualitative analysis. It not only reveals the performance differences between the current mainstream models in mathematical reasoning tasks but also verifies the key role of cue design and structured scoring mechanisms in improving the interpretability of the models.
The results show that GPT-4o has strong stability in complex tasks, especially in statistical inference and optimisation problems. Gemini has a complete chain of reasoning and clear expression, which is advantageous in structured issues. DeepSeek has gone well in computational problem solving, but suffers from logic breaks and step omissions when dealing with problems with ambiguous language or many reasoning steps and missing steps.
Although all three models are suitable for basic arithmetic and algebraic calculation, advanced and graduate-level problems still challenge their completeness of reasoning and logical coherence. This study offers empirical evidence of LLMs’ mathematical reasoning capacity based on automated grading, human evaluation, and visual examination. It gives theoretical and practical support for the choice of models in an educational context and offers possibilities for future research in multi-modal tasks, prompt strategy adaptation, and human-reinforced learning. The results of this study provide theoretical support and practical paths for the selection of models, the construction of assisted teaching systems, and the optimisation of multi-step prompting strategies in subsequent educational scenarios.
Code and Data Availability
6 Author Contributions
Ruonan Wang contributed to writing (original draft) and editing, SCoT framework, and analysis. Runxi Wang contributed to Gemini API coding and analysis, the construction of UNSW dataset, analysis, writing and editing. Chengfeng Wu contributed to DeepSeek API coding, the construction of the UNSW dataset, analysis, visualisation, and editing. Qinglin Zhou contributed to GPT-4o API coding, analysis, visualisation and design of the experiments. Yuwen Shen contributed to writing and editing, analysis and results.
R. Chandra contributed to conceptualisation, project supervision, editing, and analysis.
Appendix
6.1 Examples from mathematical problems
Figure 4: Selected questions and LLM answers from the GSM8K dataset
(a) MATH500 example 1
(b) MATH500 example 2
Figure 5: Selected examples from the MATH500 dataset